A ball is thrown vertically upward with a  velocity u from the top of a tower.if it strikes the ground with velocity 3u, the time taken by the ball to reach the ground is

Grade:11

4 Answers

Nicho priyatham
8 years ago
4u/g is the ans correct 
 
Nicho priyatham
8 years ago
well it is so simple a=(V-U)/t
here a=-g v=-3u u=u 
find t
Mahin Tazbiya
6 years ago
initial velocity = u =-u
final velocity = v = 3u
acceleration due to gravity = g
       v = u + gt
      3u = -u + gt
3u+u/g = t
4u/g = t
Anu
5 years ago
Well
Since u is going against gravity, u = -u
And final velocity is 3u so,  v = 3u 
Time can be calculated 
V = u + gt 
3u = -u + gt
So,  4u/g = t
Hope this will help. 
4u = gt
So,  t = 4u